OpenStudy (anonymous):

Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. Please help 3(3x+6) = 4(7x−1) 9x+18 = 28x−1 9x−9x+18 = 28x−9x−1 18 = 19x−1 18+1 = 19x−1+1 19 = 19x 1 = x

OpenStudy (anonymous):

4

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

im just kidding

OpenStudy (anonymous):

-.- youre hilarious.

OpenStudy (anonymous):

its 6x plus 7x minus 4x plus 2x divided 2

OpenStudy (johnweldon1993):

Well if the answer he got is x = 1.....he can check that by plugging in 1 for every 'x' he sees in the original equation and seeing if it works our correctly

OpenStudy (anonymous):

x=1 then its 6

OpenStudy (anonymous):

@johnweldon1993 is he right tho?

OpenStudy (johnweldon1993):

I'm not entirely sure what he is saying... but never just accept answers... Try taking the 1 he received and plugging it in for 'x' in the original equation... \[\large 3(3x+6) = 4(7x−1) \] becomes \[\large 3(3+6) = 4(7−1) \] \[\large 3(9) = 4(6) \] \[\large27 = 24 \] Well that isn't correct...so he is wrong

OpenStudy (johnweldon1993):

So lets figure out where he went wrong \[\large 3(3x+6) = 4(7x−1) \] \[\large 9x + 18 = 28x - 4\] ^ He went wrong there ^ \[\large 19x = 22\] \[\large x = \frac{22}{19}\]

OpenStudy (anonymous):

ahhh so that's what his answer should be right?

OpenStudy (johnweldon1993):

correct @marissa1111